M(1,2) is the midpoint of line segment RS. If S has the coordinates (8,10), find the coordinates of R
Answer by stanbon(75887)

M(1,2) is the midpoint of line segment RS. If S has the coordinates (8,10), find the coordinates of R
---
Let the coordinates of "R" be (x,y)
----
(x+8)/2 = 1
(y+10)/2 = 2
----
x+8 = 2
y+10 = 4
---
x = -6
y = -6
---------------
Ans: R(-6.-6)
=================
Cheers,
Stan H.